OpenStudy (habbababba1):

Using point-slope form, write the equation of the line that passes through the point (-4/12) and has a slope of (-3/4)

OpenStudy (jtresatti2):

no idea

OpenStudy (habbababba1):

then y answer?

OpenStudy (sheaspease):

-3x-y+0=0 (y-y1)=m(x-x1) (y-12)=-3(x--4) y=-3x-12+12 y=-3x+0 -3x-y+0=0

OpenStudy (sheaspease):

@habbababba1 does this help?

OpenStudy (habbababba1):

a little thank you

OpenStudy (sheaspease):

yw